# Mimicking Human Intuition: Cognitive Belief-Driven Q-Learning

### v1

@@ 1 Introduction
11footnotetext: Correspondence to: Xingrui Gu xingrui_gu@berkeley.edu and Hangyu Mao hy.mao@pku.edu.cn
Reinforcement learning (RL) algorithms aim to learn optimally rewarding behaviors by modeling how an agent acquires optimal strategies through a trial-and-error process within an environment (Sutton & Barto, 2018; Sutton et al., 1999). Although reinforcement learning has achieved significant success in areas like gaming, autonomous driving, and robotics, current algorithms continue to encounter challenges in addressing decision-making issues within complex, dynamic, and uncertain environments (Wu et al., 2024; McAleer et al., 2024; Xu et al., 2020; Watkins & Dayan, 1992; Silver et al., 2016; Mnih et al., 2015; Zhang et al., 2024; Mao et al., 2020; 2022; Guss et al., 2021).
@@
To overcome these limits, we attempt to solve the problem using a novel approach: Cognitive Science, often seen as a manifestation of human intuition. In this domain, humans typically construct and adjust mental models’ subjective beliefs when confronted with uncertainty to predict future events and make corresponding decisions (Peterson & Beach, 1967; Hastie & Dawes, 2009; Gigerenzer et al., 1991). These mental models, grounded in the cognition and experience of the world, empower humans to assess the potential consequences of various actions and make effective choices in complex settings. Notably, effectively managing uncertainty during decision-making is essential, as it directly influences both the efficiency of learning and the robustness of decisions (Kochenderfer, 2015). By leveraging this mechanism, we apply similar mental model theories to RL to improve the performance and adaptability of algorithms in various environments.
@@
We present a novel direction for enhancing uncertainty optimization in deep Q-learning by integrating cognitive science’s mental model with expected utility theory (Mongin, 1998). We propose Cognitive Belief-Driven Q-Learning (CBDQ), seen in Figure 1, an off-policy Deep Q-Learning algorithm applicable to both discrete and continuous states. Specifically, CBDQ incorporates:
@@
(2) Human Cognitive Clusters, implemented using the K-means algorithm (Ikotun et al., 2022), emulate how humans categorize information by grouping similar states within the environment’s state space. This method mirrors human cognition, where stimuli or situations are naturally classified into distinct categories, and serves as an efficient tool for state representation extraction. The model compresses high-dimensional data by clustering the state space into meaningful, low-dimensional representations, capturing essential environmental features and reducing learning complexity.
@@ 2.1 Advancements mathematical optimization in Q-Learning
Despite efforts to address overestimation bias, Double Q-Learning (Hasselt, 2010) only partially reduces maximization bias and may still cause underestimation in noisy environments, potentially leading to convergence to near-optimal rather than optimal solutions (Weng et al., 2020; Ren et al., 2021). Wang et al. (2021) proposed ensemble Q-learning as an alternative, using multiple Q-function approximators and conservatively selecting the minimum value. However, this strategy also risks underestimation and performance variability due to approximation errors and the limitations of a fixed ensemble size. In recent years, researchers have developed innovative Q-learning algorithms. For example, (Bas-Serrano et al., 2021) introduced Logistic Q-Learning, using a homoscedastic logistic noise model to reframe value learning via linear programming. Garg et al. (2023) proposed Extreme Q-Learning (XQL), which utilizes a Gumbel noise source along with the LINEX loss function to more effectively capture the asymmetry in Q-value distributions. Hui et al. (2023) developed Double Gumbel Q-Learning (DoubleGum), incorporating two heteroscedastic Gumbel noise sources and an adjustable pessimism factor to mitigate estimation bias. These approaches offer crucial theoretical and practical advancements for resolving Q-learning biases. While these optimization-based methods have partially addressed estimation bias, they remain incremental improvements within the Q-learning framework. Logistic Q-Learning has limited use in complex environments, XQL struggles with diverse uncertainties, and though DoubleGum offers a broader theoretical framework, it still faces key challenges, notably the lack of proven convergence. One might question: Is there a unique way of thinking that can improve algorithms like Q-learning?
@@ 2.2 Learning Process Insight Algorithms in Reinforcement Learning
Ongoing development in human-like science and RL have increasingly focused on integrating human-like reasoning and beliefs, key components of learning process-oriented algorithms. These models aim to emulate human decision-making by adapting beliefs and strategies based on experience. Complementing these efforts, Barber (2012) discusses Bayesian reasoning frameworks that incorporate prior knowledge to manage uncertainty effectively. Building on this, Carroll et al. (2019) explored collaboration by integrating learned human policies into Q-learning. More recently, Zhang et al. (2021) introduced Solipsistic Reinforcement Learning, extracting human-perspective state representations, while Hu et al. (2021) developed Off-Belief Learning (OBL), allowing agents to reason about others’ actions with dynamic beliefs. Additionally, O’Donoghue (2021) proposed Variational Bayesian Reinforcement Learning, which offers a novel approach to balancing exploration and exploitation using a risk-seeking utility function. This method introduces a new Bellman operator with associated fixed points, termed ’knowledge values,’ which compress both expected future rewards and epistemic uncertainty into a single value. These approaches enhance AI adaptability and align reinforcement learning with human cognition.
@@ 4.2 Evolving Beliefs in Q-Learning
We conducted experiments based on Example 6.7 in Sutton & Barto (2018)’s research (MBP) to verify the effectiveness of dynamically updating the subjective belief model. Four smoothing strategies, each employing a different fixed subjective belief probability model (Softmax, Clipped Max, Clipped Softmax, and Bayesian Inference), detail in Appendix C are compared with Q-learning and Double Q-learning to demonstrate the universality and accuracy of the dynamic updating mechanism for managing uncertainty.

### v2

@@ 1 Introduction
22footnotetext: Correspondence to: Xingrui Gu xingrui_gu@berkeley.edu and Hangyu Mao hy.mao@pku.edu.cn
Reinforcement learning (RL) algorithms aim to learn optimally rewarding behaviors by modeling how an agent acquires optimal strategies through a trial-and-error process within an environment (Sutton & Barto, 2018; Sutton et al., 1999). Although RL has achieved significant success in areas like gaming, autonomous driving, and robotics, current algorithms continue to encounter challenges in addressing decision-making issues within complex, dynamic, and uncertain environments (Wu et al., 2024; McAleer et al., 2024; Xu et al., 2020; Watkins & Dayan, 1992; Silver et al., 2016; Mnih et al., 2015; Mao et al., 2020; Zhang et al., 2024; Mao et al., 2022; Guss et al., 2021).
@@
To overcome these limits, we propose to use a novel approach: Cognitive Science, often seen as a manifestation of human intuition. In this domain, humans typically construct and adjust mental models’ subjective beliefs when confronted with uncertainty to predict future events and make corresponding decisions (Peterson & Beach, 1967; Hastie & Dawes, 2009; Gigerenzer et al., 1991). These mental models, grounded in the cognition and experience of the world, empower humans to assess the potential consequences of various actions and make effective choices in complex settings. Notably, effectively managing uncertainty during decision-making is essential, as it directly influences both the efficiency of learning and the robustness of decisions (Kochenderfer, 2015). By leveraging this mechanism, we apply similar mental model theories to RL to improve the performance and adaptability of algorithms in various environments.
@@
We present a novel direction for enhancing uncertainty optimization in deep Q-learning by integrating cognitive science’s mental model with expected utility theory (Mongin, 1998). We propose Cognitive Belief-Driven Q-Learning (CBDQ), shown in Figure 1, an off-policy deep Q-Learning algorithm applicable to both discrete and continuous states. Specifically, CBDQ incorporates:
@@
(2) Human Cognitive Clusters, implemented using the K-means algorithm (Ikotun et al., 2022), emulate how humans categorize information by grouping similar states within the environment’s state space. This method mirrors human cognition, where stimuli or situations are naturally classified into distinct categories, and serve as an efficient tool for state representation extraction. The model compresses high-dimensional data by clustering the state space into meaningful, low-dimensional representations, capturing essential environmental features and reducing learning complexity.
@@ 2.1 Advancements mathematical optimization in Q-Learning
Despite efforts to address overestimation bias, Double Q-Learning (Hasselt, 2010) only partially reduces maximization bias and may still cause underestimation in noisy environments, potentially leading to convergence to near-optimal rather than optimal solutions (Weng et al., 2020; Ren et al., 2021; Wang et al., 2021) proposed ensemble Q-learning as an alternative, using multiple Q-function approximators and conservatively selecting the minimum value. However, this strategy also risks underestimation and performance variability due to approximation errors and the limitations of a fixed ensemble size. In recent years, researchers have developed innovative Q-learning algorithms. For example, (Bas-Serrano et al., 2021) introduced Logistic Q-Learning, using a homoscedastic logistic noise model to reframe value learning via linear programming. (Garg et al., 2023) proposed Extreme Q-Learning (XQL), which utilizes a Gumbel noise source along with the LINEX loss function to more effectively capture the asymmetry in Q-value distributions. (Hui et al., 2023) developed Double Gumbel Q-Learning (DoubleGum), incorporating two heteroscedastic Gumbel noise sources and an adjustable pessimism factor to mitigate estimation bias. These approaches offer crucial theoretical and practical advancements for resolving Q-learning biases. While these optimization-based methods have partially addressed estimation bias, they remain incremental improvements within the Q-learning framework. Logistic Q-Learning has limited use in complex environments, XQL struggles with diverse uncertainties, and though DoubleGum offers a broader theoretical framework, it still faces key challenges, notably the lack of proven convergence. One might question: Is there a unique way of thinking that can improve algorithms like Q-learning?
@@ 2.2 Learning Process Insight Algorithms in Reinforcement Learning
Ongoing development in human-like science and RL have increasingly focused on integrating human-like reasoning and beliefs, key components of learning process-oriented algorithms. These models aim to emulate human decision-making by adapting beliefs and strategies based on experience. Complementing these efforts, Barber(Barber, 2012) discusses Bayesian reasoning frameworks that incorporate prior knowledge to manage uncertainty effectively. Building on this, Carroll(Carroll et al., 2019) explored collaboration by integrating learned human policies into Q-learning. More recently, Zhang(Zhang et al., 2021) introduced Solipsistic Reinforcement Learning, extracting human-perspective state representations, while Hu(Hu et al., 2021) developed Off-Belief Learning (OBL), allowing agents to reason about others’ actions with dynamic beliefs. Additionally, O’Donoghue(O’Donoghue, 2021) proposed Variational Bayesian Reinforcement Learning, which offers a novel approach to balancing exploration and exploitation using a risk-seeking utility function. This method introduces a new Bellman operator with associated fixed points, termed ’knowledge values,’ which compress both expected future rewards and epistemic uncertainty into a single value. These approaches enhance AI adaptability and align reinforcement learning with human cognition.
@@ 4.2 Evolving Beliefs in Q-Learning
We conducted experiments based on Example 6.7 in (Sutton & Barto, 2018)’s research (MBP) to verify the effectiveness of dynamically updating the subjective belief model. Four smoothing strategies, each employing a different fixed subjective belief probability model (Softmax, Clipped Max, Clipped Softmax, and Bayesian Inference), detail in Appendix C are compared with Q-learning and Double Q-learning to demonstrate the universality and accuracy of the dynamic updating mechanism for managing uncertainty.